OpenStudy (anonymous):

Please check my answer, thanks! . The triangles shown are similar. What is the value of x? (I'll draw it)

OpenStudy (anonymous):



OpenStudy (jdoe0001):

hmm what proportion did you use to get 10 anyway?

OpenStudy (anonymous):

4/8=5/x = x=10

OpenStudy (jdoe0001):

hmm that'd work as well, yes \(\bf \cfrac{x}{5}=\cfrac{6}{3}\implies x=\cfrac{\cancel{30}}{\cancel{3}}\implies x=10\)

OpenStudy (anonymous):

so I'm correct?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

k thanks